OpenStudy (anonymous):

How does charge travel in circuit if a capacitors is made of plates that do not contact?

OpenStudy (jamesj):

Charge doesn't cross capacitors so if you have a circuit with a capacitor in it, then charge does not make a look in the circuit through the capacitor. Hence the answer is: it doesn't! But if you have an AC source for instance, the charge flows back and forward.

OpenStudy (anonymous):

if charge cross from one plate to another capacitor is broken(disable)

OpenStudy (anonymous):

I thought that charge was a quantity but apparently its a property. definition of capacitance is C = Q/V v is voltage which is electrical potential energy per charge Can anyone make sense of these statements? More electrical potential energy per unit charge = less capacitance (ability to store charge) -- WHY Less electrical potential energy per unit charge = more capacitance (ability to store charge) -- again WHY I don't understand how increasing the distance between two charges (one fixed one not) thus decreasing electrical potential energy leads to more capacitance or why decreasing the distance between two charges (one fixed one not) thus increasing electrical potential energy leads to less capacitance can someone help clear this pleasE?

OpenStudy (anonymous):

I thought increasing the distance between two or more charges weakens their influence on each other or force and that decreasing the distance strengthens their influence on each other r force. In what way Is force between two charges proportional to capacitance?

OpenStudy (anonymous):

"I don't understand how increasing the distance between two charges (one fixed one not) thus decreasing electrical potential energy leads to more capacitance." You are not decreasing potential energy (read: voltage) across plates. Voltage is supplied by a voltage source, like a battery. Voltage across the capacitor will be the same as voltage of the battery in a circuit like this: When you increase the distance of the plates you are actually reducing capacitance so since the voltage doesn't change that must mean that the charge on the plate changes, since Q=V*C. Hope this helps

OpenStudy (jamesj):

"I thought that charge was a quantity but apparently its a property." No, charge is a quantity. In the next equation you write down: C = Q/V Q is the charge.

OpenStudy (jamesj):

You're asking quite a few questions bout capacitance that are a little bit off. I suggest going back and understanding from basics what capacitance is and how it works. Your text book is a good thing to look at and you might also enjoy this lecture: http://ocw.mit.edu/courses/physics/8-02-electricity-and-magnetism-spring-2002/video-lectures/lecture-7-capacitance-and-field-energy/